# Learning (Approximately) Equivariant Networks via Constrained Optimization

### v1

@@ 5 Experimental Setup and Results
    Theory vs. Empirics: To what extent do the theoretical equivariance-error bounds (Theorem 4.2) predict the actual deviations observed during training?
@@
Table 1: Test MSE (×$\times 10^{-2}$) on the CMU MoCap dataset for Subject #9 (Run) and Subject #35 (Walk). Results are shown for the standard equivariant models, EGNO models trained with an equality constraint ($\mathrm{ACE}\dagger$) and an EGNO model trained with an resilient inequality constraint ($\mathrm{ACE}\ddagger$). The partially equivariant EGNO ($f^{\mathrm{eq}}+f^{\mathrm{neq}}$) outperforms both the standard and equality-constrained projection $f^{\mathrm{eq}}$ models, with the adaptive-resilience variant achieving the lowest error on both subjects.
[…]
MPNN [9] 66.4$\pm$2.2 36.1$\pm$1.5
RF [62] 521.3$\pm$2.3 188.0$\pm$1.9
TFN [13] 56.6$\pm$1.7 32.0$\pm$1.8
SE(3)-Tr.[42] 61.2$\pm$2.3 31.5$\pm$2.1
EGNN [43] 50.9$\pm$0.9 28.7$\pm$1.6
EGNO (report.) [46] 33.9$\pm$1.7 8.1$\pm$1.6
EGNO (reprod.) [46] 35.3$\pm$3.2 8.5$\pm$1.0
[…]
To address RQ1, we begin with the N-Body simulations dataset [63]. In Figure 2 (left), we compare the validation MSE (mean±$\pm$±std) across 2000 epochs for the “vanilla” SEGNN and an equivariant projection $f^{\mathrm{eq}}$ obtained by training with ACE. Remarkably, $f^{\mathrm{eq}}$ trained with ACE converges more rapidly and attains a lower final MSE than the vanilla baseline, demonstrating both accelerated training dynamics and greater stability. Figure 2 (right) shows test MSE with respect to the training set size. Here, our $f^{\mathrm{eq}}$ trained on only 5000 samples matches (and slightly exceeds) the vanilla SEGNN’s performance at 9000 samples, corresponding to a $\approx 44\%$ reduction in required data. Table 2 (left) presents test MSE results (×$\times 10^{-3}$) for various SEGNN variants and baselines. Our SEGNN trained with ACE, $f^{\mathrm{eq}}$, outperforms all other baselines, including the heuristic-based relaxed equivariance scheme of Pertigkiozoglou et al. [41]  and, after projection, matches the performance of a non-equivariant SEGNN that retains the same architecture $f^{\mathrm{eq}}+f^{\mathrm{neq}}$ but, without using ACE, relies on explicit $\mathsf{SO}(3)$ augmentations to learn equivariance.
For molecular property regression, we evaluate on the QM9 dataset [64, 65] using the invariant SchNet [12] and our ACE variant. Due to its strict invariance, SchNet may severely hinder convergence without any relaxation; indeed, as shown in Table 2 (right), incorporating ACE during training reduces the mean absolute error (MAE) across all quantum-chemical targets. To assess the impact of ACE on a more flexible architecture, we compare a standard SEGNN [44] against its ACE counterpart on QM9. As shown in Table 2 (right), the ACE SEGNN lowers the MAE on most targets; however, the overall performance remains largely comparable. Collectively, these findings show that SEGNN’s $\mathsf{E(3)}$-equivariance aligns well with the QM9 dataset, facilitating optimization. However, ACE can still provide an additional boost in performance for certain molecular properties.
On the ModelNet40 classification benchmark [66], we apply ACE to the VN-DGCNN architecture [67, 45]. Figure 7 (left) shows that our ACE equivariant model $f^{\mathrm{eq}}$ achieves gains of $+2.78$ and $+1.77$ percentage points for the class and instance accuracy metrics over normal training. The partially equivariant ACE variant $f^{\mathrm{eq}}+f^{\mathrm{neq}}$ achieves comparable improvements. Both models are obtained by checkpointing the same training runs.
[…]
To address RQ2, we first evaluate convergence under input degradation on ModelNet40 [66] by randomly dropping $0\text{-}85\%$ of the points in the 3D point clouds of every shape. As shown in Figure 7 (right), the strictly equivariant baseline fails to converge under this perturbation, whereas our equality-constrained projection $f^{\mathrm{eq}}$ converges quickly to a substantially higher test accuracy. Next, we evaluate predictive performance on the CMU Motion Capture (MoCap) “Run” and “Walk” datasets [69], where we examine the effect of training with both the equality constraint and the inequality constraint of resilient constrained learning. As shown in Table 1, the equivariant projection $f^{\mathrm{eq}}$ performs similar to the vanilla EGNO. However, the partially equivariant $f^{\mathrm{eq}}+f^{\mathrm{neq}}$ outperforms both models, indicating that strictly equivariant models might not align well with the dataset. To examine whether better predictive performance can be obtained with a partially equivariant model, we perform experiments with resilient inequality constraints and observe that this training strategy further reduces test MSE on both Run and Walk compared to the purely equivariant model and the ones trained with equality constraints. These results confirm that resilience parameters can adaptively allocate non-equivariant capacity where it is most needed, improving overall performance.
To better understand these effects, we inspect how the constrained parameters evolve during training. On ModelNet40, the modulation coefficients $\gamma_{k}$ in the equality-constrained VN-DGCNN decay rapidly and reach zero within the first third of training (Figure 6, left). In the early layers, $\gamma$ rapidly decays to zero; in the deeper layers, it oscillates around zero, indicating that the optimizer initially explores non-equivariant directions before settling on a strictly equivariant solution that aligns with the data symmetries. In contrast, the same coefficients in EGNO oscillate around 0 throughout training on the MoCap dataset, suggesting that strict equivariance may be overly restrictive for this task (Figure 6, right). The learned slack values $u_{k}$ in the resilient model further support this view: as shown in Figure 5, early layers acquire larger slack, allowing flexibility where local structure dominates, while later layers remain nearly equivariant. As training progresses, the constraints tighten for all layers, increasingly penalizing non-equivariant solutions. We provide plots for all the layers in Section A.4. These dynamics demonstrate the ability of the proposed constrained learning framework to guide the model toward suitable, approximately equivariant behavior, stabilizing training under degradation, and improving performance over fully equivariant models.
Table 2: Error metrics for equality-constrained models. Left: N-body, equality-constrained SEGNN projections reach the lowest test MSE (×$\times 10^{-3}$). $\mathrm{ACE}\dagger$ represents models trained with equality constraints, $\ast$ are trained without ACE, but with $\mathsf{SO}(3)$ augmentations. Right: QM9, constraining SchNet and SEGNN reduces or matches MAE on most quantum-chemical targets.
Method MSE $\downarrow$
Linear [43] 81.9$\pm$0.00
SE(3)-Tr. [42] 24.4$\pm$0.00
RF [62] 10.4$\pm$0.00
ClofNet [70] 6.5$\pm$0.00
EGNN [43] 7.1$\pm$0.00
EGNO [46] 5.4$\pm$0.00
SEGNN [44] 5.6$\pm$0.25
@@
SchNet [12] (MAE $\downarrow$) SEGNN [44] (MAE $\downarrow$)
@@
$\mu$ 0.0511$\pm$0.0011 0.0393$\pm$0.0007 0.0371$\pm$0.0002 0.0362$\pm$0.0007
$\alpha$ 0.1236$\pm$0.0037 0.0952$\pm$0.0042 0.0991$\pm$0.0033 0.0936$\pm$0.0126
$\varepsilon_{\text{homo}}$ 0.0500$\pm$0.0006 0.0443$\pm$0.0006 0.0250$\pm$0.0005 0.0244$\pm$0.0007
$\varepsilon_{\text{lumo}}$ 0.0421$\pm$0.0005 0.0360$\pm$0.0004 0.0244$\pm$0.0003 0.0257$\pm$0.0006
$\Delta\varepsilon$ 0.0758$\pm$0.0008 0.0686$\pm$0.0011 0.0450$\pm$0.0009 0.0505$\pm$0.0091
$\langle R^{2}\rangle$ 1.4486$\pm$0.1225 0.7535$\pm$0.0426 1.2028$\pm$0.0827 1.1681$\pm$0.1890
ZPVE 0.0029$\pm$0.0003 0.0018$\pm$0.0000 0.0029$\pm$0.0004 0.0025$\pm$0.0001
$U_{0}$ 9.6429$\pm$3.2673 3.4039$\pm$0.2872 2.6163$\pm$0.4020 2.5597$\pm$0.7650
$U$ 8.3412$\pm$2.8914 3.3540$\pm$0.5006 2.5645$\pm$0.5310 2.4796$\pm$0.1440
$H$ 9.5835$\pm$4.1029 3.4403$\pm$0.4403 2.7577$\pm$0.7910 2.4569$\pm$0.5460
$G$ 8.5405$\pm$2.0311 3.6473$\pm$0.4100 2.7505$\pm$0.6970 2.5954$\pm$1.0250
$c_{v}$ 0.0379$\pm$0.0003 0.0376$\pm$0.0006 0.0389$\pm$0.0038 0.0403$\pm$0.0050
To address RQ3, we analyze how equivariance aligns with observed model behavior in the ACE resilient setting. We compute the equivariance deviation as $\mathbb{E}_{g\sim\mathsf{SO}(3)}\left[\left\|\rho_{Y}(g)f_{\theta,\gamma}(x)-f_{\theta,\gamma}\big{(}\rho_{X}(g)x\big{)}\right\|\right]$ (cf. Theorem 4.2), where the expectation is approximated using five random samples $g\in\mathsf{SO}(3)$. As shown in Figure 4, the deviation from equivariance measured for the resilient model steadily decreases and approaches values near zero by the end of training. This suggests that the equivariance constraint is effectively enforced in practice, and that partial equivariance can be preserved in a stable and interpretable way throughout training.
@@ 6 Conclusion
Appendix A Appendix / supplemental material
A.1 Proof of Theorem 4.1

### v2

@@ 5 Experimental Setup and Results
    Theory vs. Empirics: To what extent do the theoretical equivariance‐error bounds (Theorem˜4.2) predict the actual deviations observed during training?
@@
Table 1: Test MSE ($\times 10^{-2}$) on the CMU MoCap dataset for Subject #9 (Run) and Subject #35 (Walk). Results are shown for the standard equivariant models, EGNO models trained with an equality constraint ($\mathrm{ACE}\dagger$) and an EGNO model trained with an resilient inequality constraint ($\mathrm{ACE}\ddagger$). The partially equivariant EGNO ($f^{\mathrm{eq}}+f^{\mathrm{neq}}$) outperforms both the standard and equality-constrained projection $f^{\mathrm{eq}}$ models, with the adaptive-resilience variant achieving the lowest error on both subjects.
@@
[…]
5.1 Final Performance and Sample Efficiency
N-Body Simulations.
[…]
Molecular Property Regression.
We evaluate on the QM9 dataset [Rud+2012, Ram+2014] using the invariant SchNet [Sch+2017] and our ACE variant. Due to its strict invariance, SchNet may severely hinder convergence without any relaxation; indeed, as shown in Table˜2 (right), incorporating ACE during training reduces the mean absolute error (MAE) across all quantum‐chemical targets. To assess the impact of ACE on a more flexible architecture, we compare a standard SEGNN [brandstetter2022geometric] against its ACE counterpart on QM9. As shown in Table˜2 (right), the ACE SEGNN lowers the MAE on most targets; however, the overall performance remains largely comparable. Collectively, these findings show that SEGNN’s $\mathsf{E(3)}$‐equivariance aligns well with the QM9 dataset, facilitating optimization. However, ACE can still provide an additional boost in performance for certain molecular properties.
3D Shape Classification.
On the ModelNet40 classification benchmark [ModelNet], we apply ACE to the VN‐DGCNN architecture [dgcnn, deng2021vn]. Figure 7 (left) shows that our ACE equivariant model $f^{\mathrm{eq}}$ achieves gains of +2.78+2.78 and +1.77+1.77 percentage points for the class and instance accuracy metrics over normal training. The partially equivariant ACE variant $f^{\mathrm{eq}}+f^{\mathrm{neq}}$ achieves comparable improvements. Both models are obtained by checkpointing the same training runs.
[…]
5.2 Training Dynamics and Robustness
Noisy 3D Shape Classification.
To address RQ2, we first evaluate convergence under input degradation on ModelNet40 [ModelNet] by randomly dropping $0\text{-}85\%$ of the points in the 3D point clouds of every shape. As shown in Figure 7 (right), the strictly equivariant baseline fails to converge under this perturbation, whereas our equality‐constrained projection $f^{\mathrm{eq}}$ converges quickly to a substantially higher test accuracy.
Relaxed Equivariance for Motion Capture.
Next, we evaluate predictive performance on the CMU Motion Capture (MoCap) “Run” and “Walk” datasets [CMUMocap], where we examine the effect of training with both the equality constraint and the inequality constraint of resilient constrained learning. As shown in Table 1, the equivariant projection $f^{\mathrm{eq}}$ performs similar to the vanilla EGNO. However, the partially equivariant $f^{\mathrm{eq}}+f^{\mathrm{neq}}$ outperforms both models, indicating that strictly equivariant models might not align well with the dataset. To examine whether better predictive performance can be obtained with a partially equivariant model, we perform experiments with resilient inequality constraints and observe that this training strategy further reduces test MSE on both Run and Walk compared to the purely equivariant model and the ones trained with equality constraints. These results confirm that resilience parameters can adaptively allocate non‐equivariant capacity where it is most needed, improving overall performance.
Modulation Coefficients Dynamics.
To better understand the effects of ACE, we inspect how the constrained parameters evolve during training. On ModelNet40, the modulation coefficients $\gamma_{k}$ in the equality‐constrained VN-DGCNN decay rapidly and reach zero within the first third of training (Figure 6, left). In the early layers, $\gamma$ rapidly decays to zero; in the deeper layers, it oscillates around zero, indicating that the optimizer initially explores non-equivariant directions before settling on a strictly equivariant solution that aligns with the data symmetries. In contrast, the same coefficients in EGNO oscillate around 0 throughout training on the MoCap dataset, suggesting that strict equivariance may be overly restrictive for this task (Figure 6, right). The learned slack values $u_{k}$ in the resilient model further support this view: as shown in Figure 5, early layers acquire larger slack, allowing flexibility where local structure dominates, while later layers remain nearly equivariant. As training progresses, the constraints tighten for all layers, increasingly penalizing non-equivariant solutions. We provide plots for all the layers in Section˜A.8. These dynamics demonstrate the ability of the proposed constrained learning framework to guide the model toward suitable, approximately equivariant behavior, stabilizing training under degradation, and improving performance over fully equivariant models.
Table 2: Error metrics for equality-constrained models. Left: N-body, equality-constrained SEGNN projections reach the lowest test MSE ($\times 10^{-3}$). $\mathrm{ACE}\dagger$ represents models trained with equality constraints, $\ast$ are trained without ACE, but with $\mathsf{SO}(3)$ augmentations. Right: QM9, constraining SchNet and SEGNN reduces or matches MAE on most quantum-chemical targets.
Method MSE $\downarrow$
[…]
SchNet [Sch+2017] (MAE $\downarrow$) SEGNN [brandstetter2022geometric] (MAE $\downarrow$)
@@ 5.2 Training Dynamics and Robustness
[…]
5.3 Theory-Empirics Alignment
To address RQ3, we analyze how equivariance aligns with observed model behavior in the ACE resilient setting. We compute the equivariance deviation as $\mathbb{E}_{g\sim\mathsf{SO}(3)}\left[\left\|\rho_{Y}(g)f_{\theta,\gamma}(x)-f_{\theta,\gamma}\big(\rho_{X}(g)x\big)\right\|\right]$ (cf. Theorem˜4.2), where the expectation is approximated using five random samples $g\in\mathsf{SO}(3)$. As shown in Figure 4, the deviation from equivariance measured for the resilient model steadily decreases and approaches values near zero by the end of training. This suggests that the equivariance constraint is effectively enforced in practice, and that partial equivariance can be preserved in a stable and interpretable way throughout training.
6 Limitations and Further Work
While our ACE framework is broadly applicable to a range of architectures, it does incur some overhead: during training we introduce extra (non-equivariant) MLP layers, which modestly slow down training (see Appendix Table˜4). Addressing this, one avenue for future research is to “break” equivariance without adding any auxiliary parameters, e.g., by integrating constraint enforcement directly into existing layers or by designing specialized update rules that can break equivariance while preserving runtime efficiency.
Moreover, our experiments have focused on symmetry groups acting on point clouds, molecules, and motion trajectories. A natural extension is to apply our approach to other symmetry groups and data modalities, such as 2D image convolutions or 2D graph neural networks. As an initial step in this direction, our $C_{4}$-symmetric 2D convolution toy experiment (see  Section˜A.4) shows that ACE adapts appropriately, driving $\gamma$ toward zero when the target matches the imposed symmetry, and increases $\gamma$ when correcting for symmetry mis-specification, suggesting feasibility for broader 2D settings. Alongside this, co-designing network architectures that anticipate the dynamic tightening and loosening of equivariance constraints could produce models that converge even faster, require fewer samples, or require fewer extra parameters.
@@ 7 Conclusion
8 Acknowledgements
A. Manolache and M. Niepert acknowledge funding by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 – 390740016, the support by the Stuttgart Center for Simulation Science (SimTech), the International Max Planck Research School for Intelligent Systems (IMPRS-IS), and the support of the German Federal Ministry of Education and Research (BMBF) as part of InnoPhase (funding code: 02NUK078). A. Manolache acknowledges funding by the EU Horizon project ELIAS (No. 101120237). The work of L.F.O. Chamon is supported by the Agence Nationale de la Recherche (ANR) project ANR-25-CE23-3477-01 as well as a chair from Hi!PARIS, funded in part by the ANR AI Cluster 2030 and ANR-22-CMAS-0002. A. Manolache is grateful to B. Luchian for insightful discussions on symmetries and equivariant functions.
Appendix A Appendix / Supplementary Materials
A.1 Pseudo-Code for ACE
To incorporate ACE, we extend a standard equivariant model with additional parameters $\gamma_{i}$ and non-equivariant layers. The code below illustrates how to implement the architecture accordingly:
class Model($n$.Module):
def __init__(...):
self.gammas = $n$.ParameterList([
$n$.Parameter(torch.tensor(1.)) for _ in range(n_layers)
self.f_neq = $n$.ModuleList([
NoneqModule(...) for _ in range(n_layers)
self.f_eq = $n$.ModuleList([
EqModule(...) for _ in range(n_layers)
def forward(...):
for i in range(len(self.gammas)):
x = self.f_eq[i](x) + gamma[i] * self.f_neq[i](x)
Training requires introducing dual variables and a corresponding dual optimizer, while keeping the primal optimizer unchanged. The snippet below shows how to add these components and update them during training:
gammas = model.get_gammas()
dual_parameters = []
lambdas = [
torch.zeros(1, dtype=torch.float, requires_grad=False).squeeze()
for _ in gammas
dual_parameters.extend(lambdas)
# If we want to use Algorithm 2, add:
# us = [
$n$.Parameter(torch.zeros(1, requires_grad=True).squeeze())
for _ in range(len(lambdas))
# dual_parameters.extend(us)
optimizer_primal = optim.Adam/SGD/etc(model.parameters())
optimizer_dual = optim.Adam/SGD/etc(dual_parameters)
# Now, in the training routine:
loss = loss_fn(preds, y)
dual_loss = 0
slacks = [gamma for gamma in gammas]
for dual_var, slack in zip(lambdas, slacks):
dual_loss += dual_var * slack
# Explicit, but can also be obtained via autodiff
for i, slack in enumerate(slacks):
lambdas[i].grad = -slack
# If we use Algorithm 2, change to:
# loss = loss + 1/2 * torch.norm(torch.stack(us)) ** 2
# dual_loss = 0
# slacks = [torch.norm(gamma) - u for gamma, u in zip(gammas, us)]
# for dual_var, slack, u in zip(lambdas, slacks, us):
# dual_loss += (dual_var * slack) - (u * slack)
# for i, (slack, u) in enumerate(zip(slacks, us)):
# lambdas[i].grad = -slack - u
lagrangian = loss + dual_loss
lagrangian.backward()
optimizer.step()
optimizer_dual.step()
# If we use Algorithm 2, add:
# for lambda in lambdas:
# lambda.clamp_(min=0)
(a) Test MSE $(\times 10^{-3})$ on the N-Body dataset under different initializations of $\gamma$. “N/A” indicates that ACE is not used.
$\eta_{d}$ $\gamma_{\text{init}}$ Test MSE
N/A N/A $6.64\pm 0.14$
$8\times 10^{-4}$ 11 $4.97\pm 0.01$
$8\times 10^{-4}$ 0 $5.39\pm 0.33$
(b) Test MSE $(\times 10^{-3})$ on the N-Body dataset for different dual learning rates $\eta_{d}$. “N/A” indicates that ACE is not used. The primal learning rate is fixed to $\eta_{p}=9\times 10^{-4}$.
$\eta_{d}$ Test MSE
N/A $6.64\pm 0.14$
$1\times 10^{-5}$ $6.10\pm 0.25$
$8\times 10^{-5}$ $5.47\pm 0.29$
$1\times 10^{-4}$ $5.24\pm 0.29$
$8\times 10^{-4}$ $4.97\pm 0.01$
$1\times 10^{-3}$ $5.03\pm 0.09$
$8\times 10^{-3}$ $5.11\pm 0.16$
Table 3: Ablations on the N-Body dataset. (a) Effect of initializing $\gamma$. (b) Effect of varying the dual learning rate $\eta_{d}$.
A.2 Discussion regarding gamma initialization
All experiments in the main text use $\gamma_{\text{init}}=1$, as we did not find it necessary to run a separate hyperparameter search over this value. To illustrate the effect of initialization, we performed an ablation with SEGNN on the N-Body dataset where we compared $\gamma_{\text{init}}=0$ against $\gamma_{\text{init}}=1$ (Table˜3(a)).
As can be seen, both models using ACE obtain improved performance compared to removing ACE altogether ($\eta_{d}=\text{N/A}$, $\gamma_{\text{init}}=\text{N/A}$). However, initializing with $\gamma_{\text{init}}=1$ leads to the lowest test error. We observe empirically that that zero initialization encourages faster convergence in the early stages of training, but could bias the optimization toward solutions that closely adhere to the symmetry constraints, potentially limiting exploration. In contrast, initializing $\gamma_{\text{init}}=1$ gives us slower initial convergence but ultimately achieves better final accuracy.
A.3 Relationship between primal and dual learning rates
The dual optimization in ACE introduces a learning rate $\eta_{d}$ in addition to the primal learning rate $\eta_{p}$. We observe empirically that setting $\eta_{d}$ close to $\eta_{p}$ leads to stable convergence across architectures and datasets. From a practical standpoint, it is often sufficient to first tune $\eta_{p}$ and then either set $\eta_{d}=\eta_{p}$ or choose a slightly smaller value (e.g., $\eta_{d}=8{\times}10^{-4}$ when $\eta_{p}=9{\times}10^{-4}$). While default choices for $\eta_{p}$ (as in official implementations) tend to perform well, we find that $\eta_{d}$ is less sensitive and can be specified by simple proportional scaling. To quantify this effect, we conducted a sensitivity study with SEGNN on the N-Body dataset with $1000$ samples. Table 3(b) reports the test MSE (mean $\pm$ std) under different $\eta_{d}$ values, including a baseline without ACE (“N/A”). The best performance was consistently achieved when $\eta_{d}\approx\eta_{p}$, confirming that matching the dual and primal learning rates provides a robust default setting.
A.4 Correcting symmetry mis-specification with ACE
Figure 8: C4 symmetry toy experiment with ACE (resilient constraints). We want to transform a square to other shapes. The strictly equivariant baseline succeeds only when the target shares $C_{4}$ symmetry (top row). Under ACE with resilient constraints, the learned coefficient $\gamma$ adapts: in the aligned case, $\gamma$ decreases towards zero, and the outputs match; in the mismatched cases (second and third rows), $\gamma$ increases, adding a missing signal that corrects the symmetry mis-specification. For visualization, outputs are min–max normalized per panel.
Motivated by analyses of symmetry breaking in 2D CNNs [symmetrybreaking], we take a different route — we do not use relaxed group convolutions; instead, we keep the convolution strictly $C_{4}$–equivariant and relax equivariance via ACE by adding a non-equivariant branch $\gamma f^{\mathrm{neq}}$ to an equivariant $C_{4}$ CNN $f^{\mathrm{eq}}$.
[…]
A.5 Controlling the scale of the non-equivariant component.
The non-equivariant layer $f^{\text{neq},i}_{\theta}$ can in principle dominate the constraint $\gamma_{i}$ if its output norm grows too large. To prevent this, we assume that each such layer is $M$-Lipschitz and a bounded operator, i.e., $\|f^{\text{neq},i}_{\theta}(x)\|\leq B\|x\|$ for all inputs $x$ during training. This condition can be enforced by reparameterizing the weights, for example by using Spectral Normalization [miyato2018spectral], which guarantees $B\leq 1$.
From a practical standpoint, we found that enforcing exact equivariance (Algorithm˜1) did not require explicit use of Spectral Normalization: during model selection, we checkpoint based on the validation performance of the equivariant projection $f_{\theta}^{\text{eq}}$, which naturally disregards the non-equivariant part. However, in the case of Resilient Constrained Learning (Algorithm˜2), Spectral Normalization becomes important for reducing the equivariance error of $f_{\theta}^{\text{eq}}+f_{\theta}^{\text{neq}}$ (see Fig.˜4). This is a consequence of $B$ not being bounded if the weights of $f^{\text{neq},i}_{\theta}$ are not reparameterized. Hence, Spectral Normalization provides a simple and effective mechanism to control the scale of the non-equivariant branch and to reduce the equivariance error.
A.6 Proof of Theorem 4.1